National Income Accounts.
GNP = 8,800. T = 2,000. C = 5,000. G = 2,200.
X = 500. M = 900. GDP = 8,600. Dep. = 500.
Determine Foreign Lending
1
Expert's answer
2016-05-06T12:35:02-0400
GNP = 8,800. T = 2,000. C = 5,000. G = 2,200.
X = 500. M = 900. GDP = 8,600. Dep. = 500.
A country’s foreign lending is, by definition, the excess of domestic saving over domestic investment:
NFI = s*GNP - i*GNP, where NFI is net foreign investment (U.S. investment abroad net of foreign investment in the United States), GNP is Gross National Product, s = Gross Saving/GNP, and i = Gross Investment/GNP.
Gross investment I = GDP - G - C - (X - M) = 8600 - 2200 - 5000 - (500 - 900) = 1800.
In equilibrium X + G + I = M + T + S, so gross saving S = X + G + I - M - T = 500 + 2200 + 1800 - 900 - 2000 = 1600.
So, NFI = 1600/8800*8800 - 1800/8800*8800 = -200.